satellite73 (satellite73):

find sin(x) if cot(x)=-4

OpenStudy (anonymous):

Lol...@satellite73 r u alright ?

OpenStudy (anonymous):

Find sin θ if cot θ = –4 and cos θ < 0.

OpenStudy (anonymous):



OpenStudy (anonymous):

first we forget about the minus sign, we will worry about that later i just drew a right triangle with the "adjacent side" of 4 and the "opposite side" of 1 because \(\frac{4}{1}=4\)

OpenStudy (anonymous):

what is missing is the hypotenuse, and we find that by pythagoras \[h^2=1^2+4^2\] \[h^2=17\] \[h=\sqrt{17}\] now we can finish labelling the triangle

OpenStudy (amistre64):

\[(sin^2 + cos^2 = 1)/sin^2\] \[1 + cot^2 = csc^2\] \[1 + (-4)^2 = csc^2\] \[17 = csc^2\] \[\sqrt{17} = csc\] \[\frac{\sqrt{17}}{17} = sin\]

OpenStudy (amistre64):

maybe do a +- for the sqrt to account for a quadrant

OpenStudy (anonymous):

\[\frac{1}{\sqrt{17}}=\frac{1}{\sqrt{17}}\times \frac{\sqrt{17}}{\sqrt{17}}=\frac{\sqrt{17}}{17}\]